OpenStudy (anonymous):

What is the Indefinite integral of: (1)/(x*root(4(x^2)-1))

OpenStudy (anonymous):

Do you have the answer on an answer sheet? I just need to check.

OpenStudy (anonymous):

one second please

OpenStudy (anonymous):

arctan of (1/root((4x^2)-1))

OpenStudy (anonymous):

Unfortunately, I'm not as far as you into the topic of integration. Maybe someone else here can be more of an assistance to you. @hartnn

hartnn (hartnn):

i was thinking of a u-substitution 4x^2 = sec u so that root(4(x^2)-1) = root (sec^2 u-1) = tan u but i am still working on it....donno whether it'll give us required answer.

hartnn (hartnn):

***4x^2 = sec^2 u

OpenStudy (anonymous):

the u should = root(4(x^2)-1) but i do not know what to do from there to solve

hartnn (hartnn):

is it given/asked to ' u should = root(4(x^2)-1)' ?? because i don't see how that wil help...

OpenStudy (anonymous):

i checked on wolframalpha same exact integral idk :/

OpenStudy (anonymous):

thats how it solved it

hartnn (hartnn):

let me try x=(sec u)/2 dx = sec u tan u /2 sec u tan u /(2 secu /2 * (tan u)) = 1 it works very nicely..

hartnn (hartnn):

did u get that^ ?

OpenStudy (anonymous):

let me try now =)

OpenStudy (blurbendy):

you could let u = du = \[\frac{ 4x }{ \sqrt{4x^2 -1} }dx\] = \[\int\limits_{}^{} 1 / (u^2 +1) du\] = arctan(u) + C Sub back for u

hartnn (hartnn):

\(\large \int \dfrac{\sec u \tan u du }{2 (\sec u/2)(tan u)}=\int 1du=u+c\)

OpenStudy (anonymous):

blurbendy what is your u ?

OpenStudy (blurbendy):

oh sorry, u = \[\sqrt{4x^2 - 1}\]

OpenStudy (anonymous):

so how did you end up substituting and canceling how did that u get you to the form of arctan ?

OpenStudy (anonymous):

the roots end up cancelling

OpenStudy (blurbendy):

arctan is the anti-derivative of anything in the form: 1 / (x^2 + 1) Just a rule of thumb basically.

OpenStudy (anonymous):

4x doesnt equal 1

OpenStudy (anonymous):

and it is not the derivative of 4x^2